# Solving the Optimal Experiment Design Problem with Mixed-Integer Convex Methods

### v5

@@ Abstract
We tackle the Optimal Experiment Design Problem, which consists of choosing experiments to run or observations to select from a finite set to estimate the parameters of a system. The objective is to maximize some measure of information gained about the system from the observations, leading to a convex integer optimization problem. We leverage Boscia.jl, a recent algorithmic framework, which is based on a nonlinear branch-and-bound algorithm with node relaxations solved to approximate optimality using Frank-Wolfe algorithms. One particular advantage of the method is its efficient utilization of the polytope formed by the original constraints which is preserved by the method, unlike alternative methods relying on epigraph-based formulations. We assess our method against both generic and specialized convex mixed-integer approaches. Computational results highlight the performance of our proposed method, especially on large and challenging instances.
Keywords:
Mixed-Integer Non-Linear Optimization Optimal Experiment Design Frank-Wolfe Boscia
@@ 1 Introduction
$\min_{\bm{\theta}\in{\mathbb{R}}^{n}}\|A\bm{\theta}-\mathbf{y}\|,$ (1)
@@
The problem is running all experiments, potentially even multiple times to account for errors, is often not realistic because of time and cost constraints. Thus, OEDP deals with finding a subset of size $N$ of the experiments providing the “most information” about the experiment space (Pukelsheim, 2006; de Aguiar et al., 1995). The number of allowed experiments $N$ has to be larger than $n$ to allow the regression model in Equation 1 to be solved. It is, however, assumed to be less than $m$.
@@
In general, OEDP leads to a Mixed-Integer Non-Linear Problem (MINLP). There has been a lot of development in the last years in solving MINLP (Kronqvist et al., 2019). Nevertheless, the capabilities of current MINLP solvers are far away from their linear counterparts, the Mixed-Integer Problem (MIP) solvers (Bestuzheva et al., 2023), especially concerning the magnitude of the problems that can be handled. Therefore, instead of solving the actual MINLP, a continuous version of OEDP, called the Limit Problem, is often solved and the integer solution is created from the continuous solution by some form of rounding (Pukelsheim, 2006). This does not necessarily lead to optimal solutions, though, and the procedure is not always applicable to a given continuous solution either.
@@
The goal of this paper is to compare the performance of different MINLP approaches for OEDP problems. A special focus is put on our newly proposed framework Boscia (Hendrych et al., 2023) which can solve larger instances and significantly outperforms the other examined approaches. Since it leverages a formulation and solution method that differs from other approaches, we establish the convergence of the Frank-Wolfe algorithm used for the continuous relaxations on the considered OEDP problems in Section 3. The different solution methods are detailed in Section 4 and the evaluation of the computational experiments can be found in Section 5.
1.0.1 Related Work
@@
As mentioned, one established method of solution is the reduction to a simpler problem by removing the integer constraints and employing heuristics to generate an integer solution from the continuous solution. Recently, there have been more publications tackling the MINLP formulation of the Optimal Experiment Design Problem. These, however, concentrate on specific information measures, in particular the A-criterion, see (Nikolov et al., 2022; Ahipaşaoğlu, 2015), and the D-criterion, (Welch, 1982; Ponte et al., 2023; Li et al., 2024). The most general solution approach known to the authors was introduced in (Ahipaşaoğlu, 2021). It considers OEDP under matrix means which, in particular, includes the A-criterion and D-criterion. While the matrix means covers many information measures of interest, it still yields a restricting class of information functions. For example, the G-criterion and V-criterion are not included in this class of functions (de Aguiar et al., 1995). Our newly proposed framework Boscia only requires the information measures to be either $L$-smooth, i.e. the gradient is Lipschitz continuous, or generalized self-concordant, thereby covering a larger group of information functions. In addition, Boscia does not suppose any prior knowledge about the structure of the problem, being thus more flexible in terms of problem formulations. On the other hand, it is highly customizable, giving the user the ability to exploit the properties of their problems to speed up the solving process. An in-depth and unified theory for the Optimal Experiment Design Problem can be found in (Pukelsheim, 2006).
1.0.2 Contribution
Our contribution can be summarized as follows.
@@ 1.0.2 Contribution
First, we propose a unified view of multiple experiment design formulations as the optimization of a nonlinear (not necessarily Lipschitz-smooth) information function over a truncated scaled probability simplex intersected with the set of integers. Unlike most other formulations that replace the nonlinear objective with nonlinear and/or conic constraints, we preserve the original structure of the problem. Additionally, we can easily handle special cases of OEDP without any reformulations since we do not suppose a specific problem structure, unlike the approach in (Ahipaşaoğlu, 2021).
Superior solution via the Boscia framework.
@@
We use the recently proposed Boscia framework (Hendrych et al., 2023) solving MINLPs with a Frank-Wolfe method for the node relaxations of a branch-and-bound tree, and show the effectiveness of our method on instances generated with various degrees of correlation between the parameters.
1.0.3 Notation
@@
In the following let $\lambda_{i}(A)$ denote the $i$-th eigenvalue of matrix $A$; we assume that these are sorted in increasing order. Moreover, $\lambda_{\min}(A)$ and $\lambda_{\max}(A)$ denote the minimum and maximum eigenvalue of $A$, respectively. Further, let $\log\det(\cdot)$ be the log-determinant of a positive definite matrix. Given matrices $A$ and $B$ of the same dimensions, $A\circ B$ denotes their Hadamard product. Given a vector $\mathbf{x}$, $\operatorname{diag}(\mathbf{x})$ denotes the diagonal matrix with $\mathbf{x}$ on its diagonal. The cones of positive definite and positive semi-definite matrices in ${\mathbb{R}}^{n\times n}$ will be denoted by $\mathbb{S}_{++}^{n}$ and $\mathbb{S}_{+}^{n}$, respectively. We will refer to them as PD and PSD cones. For $m\in{\mathbb{Z}}_{>0}$, let $[m]=\{1,2,\dots,m\}$. Lastly, we denote matrices with capital letters, e.g. $X$, vectors with bold small letters, e.g. $\mathbf{x}$, and scalars as simple small letters, e.g. $\lambda$.
2 Optimal Experiment Design
As explained in the introduction of the paper, the matrix $A$ encodes the experiments. We are allowed to perform $N$ experiments and are interested in finding the subset with the „most" information gain. Consequently, an important question to answer is how to quantify information. To that end, we introduce the information matrix
$X(\mathbf{x})=\sum\limits_{i=1}^{m}x_{i}\mathbf{v}_{i}\mathbf{v}_{i}^{\intercal}=A^{\intercal}\operatorname{diag}(\mathbf{x})A$
@@ 2 Optimal Experiment Design
where $x_{i}\in{\mathbb{Z}}_{\geq 0}$ denotes the number of times experiment $i$ is to be performed. Throughout this paper, we will use both ways of expressing $X(\mathbf{x})$ but will favor the second representation. The inverse of the information matrix is the dispersion matrix:
@@
It is a measure of the variance of the experiment parameters (Ahipaşaoğlu, 2015). Maximizing over the information matrix is equivalent to minimizing over the dispersion matrix (Pukelsheim, 2006).
Note the following properties of the information matrix. The matrix $A^{\intercal}A$ has full rank and is positive definite. Because of the non-negativity of $\mathbf{x}$, the matrix $X(\mathbf{x})$ is in the PSD cone. In particular, $X(\mathbf{x})$ is positive definite for $\mathbf{x}\in{\mathbb{Z}}_{\geq 0}^{m}$ if the non-zero entries of $\mathbf{x}$ correspond to at least $n$ linearly independent columns of $A$. Observe that the dispersion matrix only exists if $X(\mathbf{x})$ is positive definite, i.e. has full rank. To solve the regression problem (1), the information matrix $X(\mathbf{x}^{*})$ corresponding to the optimal solution $\mathbf{x}^{*}$ of OEDP has to lie in the PD cone.
Remark 1
Experiments can be run only once or be allowed to run multiple times to account for errors. In the latter case, we will suppose upper (and lower bounds) on the number of times a given experiment can be run. The sum of the upper bounds usually greatly exceeds $N$. If non-trivial lower bounds $\mathbf{l}$ are present, their sum may not exceed $N$ otherwise there is no solution respecting the time and cost constraints.
Now we answer the question posed earlier: How can we measure information? We need a function $\phi$ receiving a positive definite matrix as input and returning a number, that is $\phi:\mathbb{S}_{++}^{n}\rightarrow{\mathbb{R}}$. We will lose information by compressing a matrix to a single number. Hence, the suitable choice of $\phi$ depends on the underlying problem. Nevertheless, there are some properties that any $\phi$ has to satisfy to qualify as an information measure.
Definition 1 (Information Function (Pukelsheim, 2006))
@@
The Loewner Ordering is an ordering on the PSD cone. Let $D,B\in\mathbb{S}_{+}^{n}$, then:
$D\succcurlyeq B\text{ if and only if }D-B\in\mathbb{S}_{+}^{n}.$
Respecting the Loewner ordering together with concavity ensures that the intuition that running more experiments should not result in less information is upheld. Scaling should not change differences in information hence we require positive homogeneity. A constant function is not useful in this setting and non-negativity is a convention. Upper semi-continuity ensures that there are no sudden jumps around the optimal objective value of the optimization problem.
The Optimal Experiment Design Problem can then be defined as:
$\begin{split}\max_{\mathbf{x}}\,&\log(\phi(X(\mathbf{x})))\\ \text{s.t. }\,&\sum_{i=1}^{m}x_{i}=N\\ &\mathbf{l}\leq\mathbf{x}\leq\mathbf{u}\\ &\mathbf{x}\in{\mathbb{Z}}^{m},\end{split}$ (OEDP)
@@
Remark 2
The $\log$ does not change the ordering and thus, is not strictly necessary. It can help reformulate the objective function, though.
A special case of non-trivial lower bounds is obtained if $n$ linearly independent experiments have non-zero lower bounds. These experiments can be summarized in the matrix $C=A^{\intercal}\operatorname{diag}(\mathbf{l})A$. Notice that $C$ is positive definite. The information matrix then becomes:
$X_{C}(\mathbf{x})=C+A^{\intercal}\operatorname{diag}(\mathbf{x}-\mathbf{l})A.$
Definition 2 (Optimal Problem and Fusion Problem)
In case the lower bounds are all zero, we call the resulting problem the Optimal Problem during this paper. Replacing the information matrix $X(\mathbf{x})$ with the fusion information matrix $X_{C}(\mathbf{x})$ in the objective in (OEDP), yields the so-called Fusion Problem.
The resulting optimization problem is an integer non-linear problem which, depending on the information function $\phi$, can be $\mathcal{NP}$-hard. The two information measures we will focus on lead to $\mathcal{NP}$-hard problems (Ahipaşaoğlu, 2021). For convenience in the latter section, we define
$\mathcal{P}:=[\mathbf{l},\mathbf{u}]\cap\left\{\mathbf{x}\in{\mathbb{R}}^{m}_{\geq 0}\,\middle|\,\sum_{i=1}^{m}x_{i}=N\right\}$ (Convex Hull)
as the feasible region of the OEDPs without the integer constraints, i.e. the convex hull of all feasible integer points. Thus, the feasible region with integer constraints will be noted by $\mathcal{P}\cap{\mathbb{Z}}^{m}$.
2.1 The A-Optimal Problem and D-Optimal Problem
The most frequently used information functions arise from the matrix means $\phi_{p}$ (Pukelsheim, 2006; Ahipaşaoğlu, 2015).
Definition 3 (Matrix Mean)
Let $B\in\mathbb{S}_{+}^{n}$ and let $\lambda(B)$ denote its eigenvalues. The matrix mean $\phi_{p}$ of $B$ is defined as
$\phi_{p}(B)=\begin{cases}\lambda_{\max}(B),&\text{for }p=\infty,\\ \left(\frac{1}{n}\operatorname{Tr}(B^{p})\right)^{\frac{1}{p}},&\text{for }p\neq 0,\pm\infty,\\ \det(B)^{\frac{1}{n}},&\text{for }p=0,\\ \lambda_{\min}(B),&\text{for }p=-\infty,\\ 0&\text{for }p=[-\infty,0]\text{ and }B\text{ singular}.\end{cases}$ (2)
Note that the matrix means function $\phi_{p}$ satisfies the requirements of Definition 1 only for $p\leq 1$ (Pukelsheim, 2006). The two most commonly used criteria arising from matrix means are the D-optimality and A-optimality criteria, corresponding to $p=0$ and $p=-1$, respectively. We convert the problems to a minimization form from this point on for homogeneity with the convention of the used solution methods.
2.1.1 D-Criterion
Choosing $p=0$ and noting that $\log\left(\det(X)^{\frac{1}{n}}\right)=\frac{1}{n}\log\det X$, yields
$\begin{split}\min\limits_{\mathbf{x}}\,&-\log\det(X(\mathbf{x}))\\ \text{s.t. }\,&\;\mathbf{x}\in\mathcal{P}\cap{\mathbb{Z}}^{m}\end{split}$ (D-Opt)

### v6

# Solving Optimal Experiment Design with Mixed-Integer Convex Methods

@@ Abstract
We tackle the Optimal Experiment Design Problem, which consists of choosing experiments to run or observations to select from a finite set to estimate the parameters of a system. The objective is to maximize some measure of information gained about the system from the observations, leading to a convex integer optimization problem. We leverage Boscia.jl, a recent algorithmic framework, which is based on a nonlinear branch-and-bound algorithm with node relaxations solved to approximate optimality using Frank-Wolfe algorithms. One particular advantage of the method is its efficient utilization of the polytope formed by the original constraints which is preserved by the method, unlike alternative methods relying on epigraph-based formulations. We assess the method against both generic and specialized convex mixed-integer approaches. Computational results highlight the performance of the proposed method, especially on large and challenging instances.
@@ 1 Introduction
$\min_{\bm{\theta}\in{\mathbb{R}}^{m}}\|A\bm{\theta}-\mathbf{y}\|,$ (1)
@@
The problem is running all experiments, potentially even multiple times to account for errors, is often not realistic because of time and cost constraints. Thus, OEDP deals with finding a subset of size $N$ of the experiments providing the ‘‘most information’’ about the experiment space (Pukelsheim, 2006; de Aguiar et al., 1995). In general, the number of allowed experiments $N$ is in the interval $[n,m]$ in order to allow a solution to the regression model in Equation 1.
@@
In general, OEDP leads to a Mixed-Integer Non-Linear Problem (MINLP). There has been a lot of development in the last years in solving MINLP (Kronqvist et al., 2019). Nevertheless, the capabilities of current MINLP solvers are far away from their linear counterparts, the Mixed-Integer Problem (MIP) solvers (Bestuzheva et al., 2021), especially concerning the magnitude of the problems that can be handled. Therefore, instead of solving the actual MINLP formulation, a continuous version of OEDP, called the Limit Problem, is often solved and the integer solution is created from the continuous solution by rounding schemes and other heuristics (Pukelsheim, 2006). This does not necessarily lead to optimal solutions, though, and the procedure is not always applicable to a given continuous solution either because it requires a certain sparsity of the continuous solution.
@@
The goal of this paper is to compare the performance of different MINLP approaches for OEDP problems. A special focus is put on the newly proposed framework Boscia (Hendrych et al., 2025b) which can solve larger instances and significantly outperforms the other examined approaches. The new method leverages Frank-Wolfe algorithms and a formulation different from the other approaches. In Section 3, we establish the convergence of the Frank-Wolfe algorithm on the continuous relaxations of the considered OEDP problems. The different solution methods are detailed in Section 4 and evaluation of the computational experiments can be found in Section 5.
1.1 Related work
@@
As mentioned, one established method of solution is the reduction to a simpler problem by removing the integer constraints and employing heuristics to generate an integer solution from the continuous solution. Recently, there have been more publications tackling the MINLP formulation of the Optimal Experiment Design Problem. These, however, concentrate on specific information measures, in particular the A-criterion, see (Nikolov et al., 2022; Ahipaşaoğlu, 2015), and the D-criterion, (Welch, 1982; Ponte et al., 2025; Li et al., 2024; Xu & Liberti, 2024; Ponte et al., 2023; Fampa & Lee, 2022; Chen et al., 2024; Ponte et al., 2024). The most general solution approach known to the authors was introduced in Ahipaşaoğlu (2021). It considers OEDP under matrix means which, in particular, includes the A-criterion and D-criterion. While the matrix means covers many information measures of interest, it still yields a restricting class of information functions. For example, the G-criterion and V-criterion are not included in this class of functions (de Aguiar et al., 1995). The newly proposed framework Boscia only requires the information measures to be either $L$-smooth, i.e. the gradient is Lipschitz continuous, or generalized self-concordant, thereby covering a larger group of information functions. In addition, Boscia does not suppose any prior knowledge about the structure of the problem, being thus more flexible in terms of problem formulations. On the other hand, it is highly customizable, giving the user the ability to exploit the properties of their problems to speed up the solving process. An in-depth and unified theory for the Optimal Experiment Design Problem can be found in Pukelsheim (2006).
1.2 Contribution
Our contribution can be summarized as follows
@@ 1.2 Contribution
First, we propose a unified view of multiple experiment design formulations as the optimization of a nonlinear (not necessarily Lipschitz-smooth) information function over a truncated scaled probability simplex intersected with the set of integers. Unlike most other formulations that replace the nonlinear objective with nonlinear and/or conic constraints, we preserve the original structure of the problem. Additionally, we can easily handle special cases of OEDP without any reformulations and additional constraints on OEDP since we do not suppose a specific problem structure, unlike the approach in Ahipaşaoğlu (2021).
Convergence analysis for BPCG.
Second, we establish convergence guarantees for Frank-Wolfe algorithms on continuous relaxations of these problems by exploiting smoothness or self-concordance, and sharpness. This is necessary in order to solve the various experimental design problems with Boscia.
Superior solution via Boscia framework.
@@
Finally, we use the recently proposed Boscia (Hendrych et al., 2025b) framework to solve the integer optimization problem with a Frank-Wolfe method for the node relaxations of the Branch-and-Bound tree and show the effectiveness of the method on instances generated with various degrees of correlation between the parameters.
1.3 Notation
@@
In the following let $\lambda_{i}(A)$ denote the $i$-th eigenvalue of matrix $A$; we assume that these are sorted in increasing order. Moreover, $\lambda_{\min}(A)$ and $\lambda_{\max}(A)$ denote the minimum and maximum eigenvalue of $A$, respectively. We define and denote the condition number of a positive semi-definite matrix $A$ as $\kappa(A)=\frac{\lambda_{\max}(A)}{\lambda_{\min}(A)}$. Further, let $\log\det(\cdot)$ be the log-determinant of a positive definite matrix. Given matrices $A$ and $B$ of same dimensions, $A\circ B$ denotes their Hadamard product. Given a vector $\mathbf{x}$, $\operatorname{diag}(\mathbf{x})$ denotes the diagonal matrix with $\mathbf{x}$ on its diagonal and let $\operatorname{Diag}(A)$ denote the diagonal of the matrix $A$. The cones of positive definite and positive semi-definite matrices in ${\mathbb{R}}^{n\times n}$ will be denoted by $\mathbb{S}_{++}^{n}$ and $\mathbb{S}_{+}^{n}$, respectively. We will refer to them as PD and PSD cones. The set of symmetric $n\times n$ matrices will be denoted by $\mathbb{S}^{n}$. Let ${\mathbb{N}}_{0}$ denote the set of positive integers including 0 and for $m\in{\mathbb{N}}$ let $[m]=\{1,2,\dots,m\}$. Lastly, we denote matrices with capital letters, e.g. $X$, vectors with bold small letters, e.g. $\mathbf{x}$, and simple small letters are scalars, e.g. $\lambda$.
2 The optimal experiment design problem
As stated in the introduction, OEDP aims to pick the subset of predefined size $N$ yielding the most information about the system. So first, we have to answer the question of how to quantify information. To that end, we introduce the information matrix
$X(\mathbf{x})=\sum\limits_{i=1}^{m}x_{i}\mathbf{v}_{i}\mathbf{v}_{i}^{\intercal}=A^{\intercal}\operatorname{diag}(\mathbf{x})A$ (2)
@@ 2 The optimal experiment design problem
where $x_{i}\in{\mathbb{N}}_{0}$ denotes the number of times experiment $i$ is to be performed. We call $\mathbf{x}\in{\mathbb{N}}_{0}^{m}$ a design. Throughout this paper, we will use both ways of expressing $X(\mathbf{x})$ but will favor the second representation. The inverse of the information matrix is the dispersion matrix
@@
It is a measure of the variance of the experiment parameters (Ahipaşaoğlu, 2015). Our goal is to obtain the ‘‘most’’ information, i.e. we want to maximize some measure of the information matrix. This is equivalent to minimizing over the dispersion matrix (Pukelsheim, 2006), thereby minimizing the variance of the parameters.
[…]
Remark.
Experiments can be run only once or be allowed to run multiple times to account for measurement errors. In the latter case, we will suppose non-trivial upper (and lower bounds) on the number of times a given experiment can be run. The sum of the upper bounds significantly exceeds $N$.
If non-trivial lower bounds $\mathbf{l}$ are present, their sum may not exceed $N$ otherwise there is no solution respecting the time and cost constraints. A special case of non-trivial lower bounds is obtained if $n$ linearly independent experiments have non-zero lower bounds. These experiments can be summarized in the matrix $C=A^{\intercal}\operatorname{diag}(\mathbf{l})A$. Notice that $C$ is positive definite. The information matrix then becomes
$X_{C}(\mathbf{x})=C+A^{\intercal}\operatorname{diag}(\mathbf{x}-\mathbf{l})A.$ (3)
We will refer to the resulting OEDP as the Fusion Experiment Design Problem or short as the Fusion Problem. The OEDP with $\mathbf{l}=0$ will be referred to as Optimal Problem. These are the two types of problems we are focusing on in this paper.
Let us now answer the question posed earlier: How can we measure information? We need a function $\phi$ receiving a positive definite matrix [^2] As previously stated, the information matrix of the optimal solution $X(\mathbf{x}^{*})$ has to have full column rank. Thus it suffices to define $\phi$ on the PD cone. as input and returning a number, that is $\phi:\mathbb{S}_{++}^{n}\rightarrow{\mathbb{R}}$. We will lose information by compressing a matrix to a single number. Hence, the suitable choice of $\phi$ depends on the underlying problem. Nevertheless, there are some properties that any $\phi$ has to satisfy to qualify as an information measure.
Definition 2.1 (Information Function (Pukelsheim, 2006)).
@@
Respecting Loewner ordering
Let $B,D\in\mathbb{S}_{+}^{n}$. We say $D\succcurlyeq B$ if and only if $D-B\succcurlyeq 0$ which is equivalent to $D-B\in\mathbb{S}_{+}^{n}$. The order thus imposed on $\mathbb{S}_{+}^{n}$ is called the Loewner Ordering. A map $\phi$ respects the Loewner ordering iff $C\succcurlyeq B\Rightarrow\phi(C)\geq\phi(B)\;\forall C,B\in\mathbb{S}_{+}^{n}$. In simple terms, carrying out more experiments should not result in information loss.
Concavity
The condition for concavity is
$\phi(\gamma B+(1-\gamma)C)\geq\gamma\phi(B)+(1-\gamma)\phi(C)\quad\forall\gamma\in[0,1]\;\forall B,C\in\mathbb{S}_{++}^{n}.$
The intuition behind this is that we should not be able to achieve the best result by performing two sets of experiments and interpolating the result. There should be a design unifying the two sets achieving a better result.
Positive homogeneity
For $\gamma>0$, we have $\phi(\gamma B)=\gamma\phi(B)$, i.e. scaling has no influence on the overall information value. It has added advantage that later in the optimization, we may forgo any factors since they do not interfere with the ordering created by $\phi$.
Nonnegativity
By convention, we have $\phi(B)\geq 0$ for all $B\in\mathbb{S}_{+}^{n}$.
Upper semi-continuity
The upper level sets $\left\{M\in\mathbb{S}_{++}\,|\,\phi(M)\geq\lambda\right\}$ are closed for all $\lambda\in{\mathbb{R}}$. This is to ensure "good" behavior at the boundary. The functions considered in this paper are all continuous.
The most frequently used information functions arise from matrix means $\phi_{p}$ (Pukelsheim (2006), Ahipaşaoğlu (2015)). The matrix means are defined as follows:
Definition 2.2 (Matrix mean).
Let $C\in\mathbb{S}_{++}^{n}$ and let $\lambda(C)$ denote the eigenvalues of $C$. The matrix mean $\phi_{p}$ of $C$ is defined as
$\phi_{p}(C)=\begin{cases}\max\lambda(C),&\text{for }p=\infty\\ \left(\frac{1}{n}\operatorname{Tr}(C^{p})\right)^{\frac{1}{p}},&\text{for }p\neq 0,\pm\infty\\ \det(C)^{\frac{1}{n}},&\text{for }p=0\\ \min\lambda(C),&\text{for }p=-\infty.\end{cases}$ (4)
If $C$ is semi-definite with $\operatorname{rank}(C)<n$, the matrix mean is defined as:
$\phi_{p}(C)=\begin{cases}\max\lambda(C),&\text{for }p=\infty\\ \left(\frac{1}{n}\operatorname{Tr}(C^{p})\right)^{\frac{1}{p}},&\text{for }p=(0,\infty)\\ 0,&\text{for }p=[0,-\infty].\end{cases}$ (5)
Note that the function $\phi_{p}$ satisfies the requirements of Definition 2.1 only for $p\leq 1$ (Pukelsheim, 2006). The different values of $p$ lead to different so-called criteria. The Optimal Experiment Design Problem using a matrix means function is then defined as
$\max_{\mathbf{x}}\,$ $\log(\phi_{p}(X(\mathbf{x})))$ (OEDP)
s.t. $\sum_{i=1}^{m}x_{i}=N$
$\mathbf{l}\leq\mathbf{x}\leq\mathbf{u}$
$\mathbf{x}\in{\mathbb{N}}_{0}^{m},$
@@
Remark.
Using the logarithm formulation, one can recover Fenchel duality results (Sagnol, 2010, Chapter 3). It can also by beneficial from a information theory point of view. Note, however, that the $\log$ does not necessarily preserve concavity.
The resulting problem (OEDP) is an integer optimization problem which, depending on the information function $\phi_{p}$, can be $\mathcal{NP}$-hard. The two information measures we will focus on lead to $\mathcal{NP}$-hard problems (Welch, 1982; Nikolov et al., 2022; Li, 2025; Li et al., 2024). Thus, these OEDP are hard to solve. Often, therefore, one solves the so-called Limit Problem obtained by letting the number of allowed experiments $N$ go to infinity. The resulting optimization problem is continuous.
$\max\limits_{\mathbf{w}}\,$ $\log(\phi_{p}(X(\mathbf{w})))$ (OEDP-Limit)
s.t. $\sum_{i=1}^{m}w_{i}=1$
$\mathbf{w}\in[0,1]^{m}$
The variable vector $\mathbf{w}$ can be interpreted as a probability distribution. Under suitable assumptions, one can generate a finite experiment design from the solution of the limit problem, see Ahipaşaoğlu (2015), Pukelsheim (2006, Chapter 12). One assumption is that the support of the solution of the Limit Problem (OEDP-Limit) is smaller than the allowed number of experiments $N$. This is often not the case (Pukelsheim, 2006). If an integer solution can be obtained, it is feasible but not necessarily optimal. Optimality can be achieved in some special cases, see Pukelsheim (2006). However, these have very specific requirements that are often not achievable.
2.1 The A-, D- and GTI-optimal experiment design problem
For this paper, we focus on two particular criteria arising from the matrix means. The two most commonly used criteria are the D-optimality and A-optimality criterion, $p=0$ and $p=-1$, respectively.
Many methods for solving MINLP expect minimization formulations. Hence, we will reformulate the problems as minimization problems by flipping the sign of the objective.
D-Criterion
Choosing $p=0$ in Definition 2.2, yields for the objective of (OEDP)
$\max\limits_{\mathbf{x}\in{\mathbb{N}}_{0}^{m}}\,\log\det\left((X(\mathbf{x}))^{\frac{1}{n}}\right)$
Note that $\log(\det(X)^{\frac{1}{n}})=\frac{1}{n}\log\det X$ and thus we can state the D-Optimal Experiment Design Problem as:
$\begin{split}\min\limits_{\mathbf{x}}\,&-\log(\det(X(\mathbf{x})))\\ \text{s.t. }\,&\sum_{i=1}^{m}x_{i}=N\\ &\mathbf{l}\leq\mathbf{x}\leq\mathbf{u}\\ &\mathbf{x}\in{\mathbb{Z}}\end{split}$ (D-Opt)